Home‎ > ‎Reports‎ > ‎Crew099‎ > ‎

day02


Crew 99 Log Book for January 30, 2011

Photos of the day





Commander's Check-In Report

Donna Viola Reporting

Crew Physical Status: good

Time Departed/Returned from EVA: EVA #1 2:10pm-3:15pm; EVA #2 4pm-5pm

Brief Narrative of Field Mission Results: practiced suiting up for EVA, learned how EVA suits affect ability to see, move, and communicate

EVA Data/Interpretations: as per EVA reports

Engineering/Hab Maintenance: n/a

Report Transmission Schedule: Chef's, EVA 1, EVA 2, Engineering, and Commander's, by 8pm

Plans for Tomorrow: sample collection EVA, practice using EVA Mission Planner

Inventory: n/a

Support Requested: Inspected EVA suits - found the following issues, please advise where necessary:
- large hole in helmet 2 faceplate
- one of the hoses from backpack 5 was connected backwards, and had gotten severed from the connector at one end.
- backpack 3 hose connection was too large to connect to helmet 3
- 1 headset was broken - could talk to other team members, but couldn't hear any communications
- had trouble using vox on radios; resorted to using push-to-talk for second EVA

Miscellaneous: packaged freeze-dried chicken and beef for subsequent crews; stored in loft

Commander's Report

Donna Viola Reporting

Today, the crew awoke to their first morning on Mars, ready to begin exploring. The ladies woke up early to practice riding the ATVs and to explore one of the commander's favorite geology sites, gathering some really cool rocks and petrified wood to show off to friends and family back home - and discovered, much to their delight, tumbleweed (something non-desert-dwellers don't get to see very often!).


Afterward, the guys ventured north to practice riding the ATVs prior to entering sim. Once they returned, the team officially started the simulation, ready to embrace their role as Marsonauts.


Over lunch, the team came to the conclusion that MDRS would be a great place to wait out a zombie apocalypse - so if that happens within the next two weeks, we should be safe (no such promises for our friends on the Mission Support team back on Earth).


The afternoon consisted of practice EVAs to learn how to operate while wearing EVA suits. Max, Heidi, and Andie went to explore Candor Chasma, looking for sites for future geologic study. Max just couldn't get enough EVAs, so he joined Kevin, Sukrit, and myself to track down the MDRS geocache - after a fairly strenuous hike, mission success!


The away team returned to the Hab triumphant, and settled in for the evening to relax and fill out some lovely surveys (not only our food study questionnaires, but also the surveys for a remote immunosuppression study being conducted by a fellow NASA Ames Academi alum).


Tomorrow, we are planning on conducting our first science EVAs - and are very excited to spend our first full day on "Mars"!


Engineering Report

Kevin Newman Reporting

Generator/Electricity:


Kitty (Diesel Generator) (Used? Yes/No): Yes


Honey (Gas Generator) (Used? Yes/No): No


Battery Charge Level:

10.62 V everstart

12.39 V armor plate

14.56 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Oil leaking from dipstick tube on Kitty, attributed to dipstick not pushed all the way in the day before. Oil cleaned up, dipstick inserted properly, no leakage since.


Honey started up as a test (started up successfully with two yanks on pull cord).


Fuel Status:


Diesel (Full/0.75/0.5/0.25/Empty):1/3


Propane Tank (Full/0.75/0.5/0.25/Empty):24%


Gasoline Tank (Full/0.75/0.5/0.25/Empty): Horizontal raised drum EMPTY. Three 5-gallon containers full.


Notes/Comments:


Oil:


10 quarts of 5w-20 on hand


1 gallon of 15w-40


Kitty Oil Quantity (# of Quarts in storage): Leave it to DG


Honey Oil Quantity (# of Quarts in storage): Leave it to DG


ATV Oil Quantity (# of Quarts in storage): 2 quarts on the engineering work bench


Car Oil Quantity (# of Quarts in storage):


Water Status:


Outside Potable Water Tank Level (inches from bottom): 18.5


Trailer Potable Water Tank Level (inches from bottom): 1/5


Hab Potable Water Tank Level (inches from bottom): 9


Potable Water Meter Reading: 414821


Notes/Comments:


GreenHab:


Condition of plants in Tank 1: Nothing


Condition of plants in Tank 2: Something green growing!


Condition of plants in Tank 3: Good


Greenhab notes/ comments:


Would you like me to send pictures of the tanks?


Observatory:


Telescope Used? (Yes/no) No


Observatory Notes:


Transportation:


Hab Car (NOT RENTAL)

Used (yes/no):No

Oil Checked (yes/no): No

Tire Status: Not checked

Odometer at end of day: 123596.3

Notes/Comments on Hab Car: Washer fluid reservoir filled


ATV 1 (Opportunity):

Used: Yes

Oil Checked: Yes

Fuel Consumed: Replaced

Tire Status: Checked, filled

Comments:


ATV 2 (Spirit):

Used: Yes

Oil Checked: Yes

Fuel Consumed: Replaced

Tire Status: Checked, filled

Comments:


ATV 3 (Viking I):

Used: No

Oil Checked: No

Fuel Consumed: None

Tire Status: Not Checked

Comments:


ATV 4 (Viking II):

Used: Yes

Oil Checked: Yes

Fuel Consumed: Replaced

Tire Status: Checked, Filled

Comments:


Heating and Ventilation:


Thermostat setting upstairs (Farenheit): 70

Actual temperature upstairs (Farenheit): 77


Thermostat setting downstairs (Farenheit): 65

Actual temperature downstairs (Farenheit): 72


Computers/Networking Infrastructure:


Notes/comments on internet/computers?: All good


General Engineering

1. Notes/Comments:

It is very helpful to have two engineers here! Max is very helpful.


There is no longer access to beneath the hab. Should I attempt to check heating strips on pipes by some other means?


EVA 1 Report

Andie Gilkey Reporting

Time: 14:11-15:15

Crew Members: Andie Gilkey, Heidi Beemer, Max Fagin

Site Location: UTM: 520000 4251000

Transit Mode: Walking

Objectives: Gain experience with the EVA suits, locate Candor Chasma using Mission Planner and GPS

Results: Experience gained w/using the Mission Planner and w/the EVA suits.

Lessons Learned: You can do a lot w/the suits, however they limit vision and mobility.

EVA 2 Report

Sukrit Ranjan Reporting

Time: 16.00-17:00


Crew Members: Sukrit Ranjan, Max Fagin, Donna Viola, Kevin Newman


Site Location: (lat, lon)=(38 deg. 24.368', -110 deg. 47.741')


Transit Mode: Walking


Objectives: Gain experience with the EVA suits, locate the MDRS geocache, sign in and leave a team LAMBDA pin.


Results: Geocache located, pin left, sign-in performed, experience gained.


Lessons Learned: You can do a lot with the suits. However, some tasks are more difficult than you might expect. For example, getting the pin out of Kevin's pocket was the hardest part of the EVA.


Cook's Report

Sukrit Ranjan Reporting

1a) Today was

-A non-cooking day.


1b) Was today a special day celebrated at the main meal with special food or activity?

- No


2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

- Leonardo da Fettuccine (Gourmet Reserves)


3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded) (Example: 2 servings potato soup stored; 2 servings rice burned/inedible, discarded)

- All prepared food was consumed.


4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

- None


5) Comments/questions for Kim and Jean, the food study investigators

- Food was quite good. Most of us liked it.